OpenStudy (anonymous):

76.Winning percentage. In a basketball league, the Falcons won 15 of their first 20 games. In order to win 60% of the total number of games, how many more games will they have to play, assuming they win only half of the remaining games? Let's let x be the total number of games that need to be played. Then what expression can you use to represent how many remaining games there would be? Then, we want 75% of the 20 + 50% of the remaining games = 60% of the total.

OpenStudy (gw2011):

I gave a response in your earlier posting. You can take a look there to see if this is the same question. If my memory serves me correctly, I believe the remaining number of games needed is 30. Adding 20 to 30 = 50 total game to be played Hallf of 30 = 15 which is assuming half of the remaining games will be won Therefore, you get: 15 previous wins + 15 wins for remaining games = 30 games won 20 previous games played + 30 additional games to be played = 50 total games Now: 30/50 = 0.60 = 60%

OpenStudy (anonymous):

ok i came up with this,hey will have to play 26 more games. making a total of 46 games played. 46 games minus the already played 20 is 26. 50% of 26 is 13, Add the 13 to the 15 already won and you get 28. 28 divided by 46 equals 60%

OpenStudy (gw2011):

Your answer comes out to 60.9% which is approximately 61% because what you got will be as follows: 28/46=0.609=60.9%. If you take my answer of 30/50, you get exactly 60%

OpenStudy (anonymous):

yea thats true..okay but with the setup that i have to use how can i write this out Let's let x be the total number of games that need to be played. Then what expression can you use to represent how many remaining games there would be? Then, we want 75% of the 20 + 50% of the remaining games = 60% of the total.

OpenStudy (anonymous):

0.75 * 20 + 0.5x = 0.6(20 + x) 15 + 0.5x = 12 + 0.6x 15 - 12 = 0.6x - 0.5x 3 = 0.1x x = 3/0.1 = 30 games left to be played.

OpenStudy (gw2011):

You're right.

OpenStudy (anonymous):

k ty

OpenStudy (gw2011):

You're welcome